5. A car moving with a speed of 90 km/hour starts to slow down and in 0.5 minutes comes to rest. Find acceleration of the car.
Answer #1

(5) Initial speed of the car = 90 km/hour

Convert this in m/sec.

90 km/hour = (90 x 1000) / (60 x 60) = 25.0 m/s

Suppose, u = 25.0 m/s

When the car stops, its speed = 0

Suppose, v = 0

Time taken by the car to come at stop, t = 0.5 min = 0.5 x 60 = 30.0 s

Suppose, a = acceleration of the car

Use the expression –

v = u + a*t

put the values-

0 = 25.0 m/s + a*30.0 s

=> a = -25.0 / 30 = -0.83 m/s^2

So, acceleration of the car = 0.83 m/s^2.

The negative sign of acceleration shows that the car is slowing down.